# Uncovering Neuroimaging Biomarkers of Brain Tumor Surgery with AI-Driven Methods

### v1

@@ Abstract
Brain tumor resection is a complex procedure with significant implications for patient survival and quality of life. Predictions of patient outcomes provide clinicians and patients the opportunity to select the most suitable onco-functional balance. In this study, global features derived from structural magnetic resonance imaging in a clinical dataset of 49 pre- and post-surgery patients identified potential biomarkers associated with survival outcomes. We propose a framework that integrates Explainable AI (XAI) with neuroimaging-based feature engineering for survival assessment, offering guidance for surgical decision-making. In this study, we introduce a global explanation optimizer that refines survival-related feature attribution in deep learning models, enhancing interpretability and reliability. Our findings suggest that survival is influenced by alterations in regions associated with cognitive and sensory functions, indicating the importance of preserving areas involved in decision-making and emotional regulation during surgery to improve outcomes. The global explanation optimizer improves both fidelity and comprehensibility of explanations compared to state-of-the-art XAI methods. It effectively identifies survival-related variability, underscoring its relevance in precision medicine for brain tumor treatment.
@@ 1 Introduction
Gliomas, the most frequent primary brain tumors, vary in aggressiveness, prognosis, and histopathology. Their treatment often involves surgical resection, followed by radiotherapy and chemotherapy. The extent of resection significantly affects survival, with surgery needing to balance tumor removal and brain function preservation [1]; known as onco-functional balance. Post-surgical brain reorganization plays a key role in recovery, but the mechanisms behind these changes are not fully understood [2]. Accurate assessment of these changes is crucial for improving patient outcomes and rehabilitation strategies.
Structural Magnetic Resonance Imaging (sMRI) may provide insights into brain reorganization, but its high dimensionality poses challenges for analysis. Machine learning (ML) techniques have demostrated superior performance over traditional methods in predicting postoperative complications and inpatient length of stay for patients with brain tumours [3, 4, 5]. They have been effectively applied to tumor segmentation, classification, and pre- and post-operative MRI analysis [6, 7, 8, 9], as well as in predicting neurosurgical outcomes, including survival and recurrence [10, 11]. The application of ML-based latent space representations, such as Principal Component Analysis (PCA) [12], has proven valuable in neuroimaging analysis by aiding in the interpretation of structural variations. However, no studies have employed latent spaces to assess structural changes in brain tumors before and after surgery, which this work addresses.
EXplainable Artificial Intelligence (XAI) techniques have effectively identified patterns in hypothesis-driven research [13, 14, 15]. Local XAI methods provide interpretations of individual model predictions, whereas global methods offer cohort-level insights into the model’s overall decision-making process, thereby enhancing our understanding of its behavior across populations. By combining these methodologies, we can improve our understanding of brain tumor-related structural reorganization and facilitate personalized treatment approaches. However, different local techniques lead to varying explanations, increasing uncertainty rather than trust [16] in AI-driven solutions. The same problem is observed in global explanations. Part of this work aims to address the problem of inter-method variability in global explanations by proposing a global XAI optimizer.
This study presents a novel framework combining latent-PCA spaces with XAI to analyze variability in brain tumors among patients who survived and those who did not, following the surgery. We introduce a global explanation optimizer to improve the accuracy and clarity of survival-related biomarkers, eliminating the inter-method variability in explanations. Additionally, we investigate the relationship between significant brain regions and survival outcomes to provide guidance that may reduce surgical fatality rates. Given the limited availability of real-world longitudinal data in neuro-oncology, particularly covering both pre- and post-surgical stages, data used in this study offers a rare opportunity to investigate the impact of surgery on brain function and survival. We focus on two factors: time (pre- and post-surgery) and survival (longer-term and shorter-term). The key contributions of our study are: (i) a global explanation optimizer to identify survival-related biomarkers, enhancing both optimal accuracy and comprehensibility of the explanations, (ii) a framework that combines XAI with neuroimaging-based feature engineering to offer novel insights into the relationship between sMRI features and survival risk, and (iii) providing suggestions and potential guidance for surgical decision-making to improve survival outcomes for patients with brain tumors.
2 The proposed framework
@@
This work combines latent space feature engineering and XAI methods to identify biomarkers related to surgical outcomes. A summary of the implemented framework is presented in Figure 1. First, sMRI images are processed, including spatial alignment, skull stripping, and masking of tumor regions to ensure consistency across subjects and facilitate subsequent analyses. The main part consists of two phases: feature engineering through dimensionality reduction and a global explanation optimizer integrated with DL networks and XAI methods.
@@ 2.1 Phase I: Feature engineering based on dimensionality reduction
We utilized PCA to extract the most relevant patterns of variation across the brain and tumor cohorts, based on four groups derived from the time/survival factors. Two main approaches were used: first, analyzing PCA component variability across groups, and second, quantifying variability across PCA components (see Figure 1):
@@ 2.1.2 Variability quantification across PCA Components
2.2 Phase II: Feature identification based on cohort-level explanations, integrated with deep learning networks and tailored to survival classification
Figure 1 presents the explainable AI framework used to identify global (cohort-level) patterns in survival outcomes after brain tumor surgery. An unsupervised learning approach was used, in which an encoder-decoder architecture learns the distribution of sMRI data. A binary classification model was then trained and validated to distinguish between subjects with shorter and longer survival. The classification task explored various combinations of frozen and fine-tuned encoder layers and an ablation study of different networks for prediction from the unsupervised learning stage to enhance generalization and improve accuracy. Finally, cohort-level explanations were identified to the survival classification task, incorporating our proposed global explanation optimizer to improve the clarity and consistency of the global explanation patterns.
@@ 2.2.1 Unsupervised learning of structural MRI
Two deep learning architectures: a convolutional autoencoder (AE) comprising three encoder and three decoder blocks, and the Swin-Unet [17] were trained in an unsupervised setting to reconstruct entire pre- and post-operative 3D structural MRI scans. An ablation study was conducted to compare the effect of two different cohort training strategies aiming to determine which approach enhances the networks’ reconstruction performance and generalization capability. A detailed explanation is given in Section 3.2.
@@ 2.2.4 The proposed global explanation optimizer of structural MRI
We extracted the first three principal components via PCA from the total cohort saliency maps, generated using three of the six widely used attribution methods employed in this study: Guided Backpropagation, Guided Grad-CAM, and Gradient SHAP. These components, along with their weighted average—computed according to the procedure described in [16]—were used as four inputs to the proposed global explanation optimizer.
@@
3 Experimental setup
@@ 3.1 Dataset
The main dataset was from Addenbrooke’s Hospital (Cambridge, UK). The UK’s Research Authority provided ethical approval (ref:19/WM/0152) and data were anonymized before analyses. The dataset consists of 49 MRI T2-weighted scans acquired both before and after surgical resection of the tumour. These scans were spatially normalized to MNI space using SPM12 (fil.ion.ucl.ac.uk/spm/) and resampled to a 1$\times$1$\times$1 mm3 resolution resulting in final image dimensions of 157$\times$189$\times$156 mm. Skull-stripping was performed using HD-BET [26]. Patients were categorized into two outcome groups: longer-term (32) and shorter-term (17) survivals. Most patients (42, $85\%$) had a glioblastoma, but there were also cases of astrocytoma (1), gliosarcoma (3) and others (3). The shorter-term survival group comprised patients who had died within 10 months after the postoperative scan. In contrast, the longer-term group included those who survived for more than 10 months. The UK’s Research Authority provided ethical approval (ref:19/WM/0152); data were anonymized before analyses
In Phase II, an additional dataset from the 2025 Brain Tumor Segmentation (BraTS) Glioma Challenge [27] was employed. Hereafter, we refer to this dataset as BraTS2025. This dataset comprises pre- and post-treatment T2-weighted MRI scans. We used a total of 1453 images (1251 pre-treatment and 202 post-treatment). These scans were used to train the unsupervised learning models (see 2.2.1). Demographic information was not provided for this dataset.
@@ 3.2 Implementation details
For survival classification, sparse categorical cross-entropy was used as the loss function, optimized with Adam. The learning rate was constant for the first 100 epochs and then reduced by a factor of 0.1 every 100 epochs. Early stopping was applied after 100 epochs of no improvement (max 400 epochs). A 5-fold CV was used. Both tasks employed data augmentation, including rotation ($[-15^{\circ},15^{\circ}]$), width/height shift (up to 20 pixels), and intensity shift (up to 20%). Hyperparameter tuning tested learning rates: $5\times 10^{-2}$, $5\times 10^{-3}$, $5\times 10^{-4}$, and $5\times 10^{-5}$ (see Figure 2a.). The XAI task used the Adam optimizer, but no data augmentation. The cost function was (4). For 3D tasks, training lasted up to 100 epochs, with early stopping after 10 epochs of no improvement beyond the first 50. Hyperparameter tuning tested the same learning rates as previously and various combinations of the $l_{1}$, $l_{2}$, and $l_{3}$ parameters in (4) with the best combination of parameters determined as $l_{1}=0.4$, $l_{2}=0.3$, $l_{3}=0.3$ and a learning rate $5\times 10^{-5}$ (see Figure  2b.). Codes were implemented in Python using PyTorch and trained on one A100 GPU with 64 GB RAM. It will be publicly available on GitHub.
@@ 3.3.2 Sparseness Metric
where $\left|\boldsymbol{\phi}^{(k)}\right|=\left(|\phi^{(k)}_{1}|,|\phi^{(k)}_{2}|,\ldots,|\phi^{(k)}_{d}|\right)$. Higher values indicate greater sparseness. A value of 1 implies that the attribution is entirely concentrated on a single feature, while 0 corresponds to equal attribution across all features.
@@ 4.1 Structural patterns identified using feature engineering based on PCA
Once the PCA components (brain: 8, tumor: 12) were computed across groups, structural variability was quantified to explore spatial differences in tumor and brain patterns. The localization of the first PCA component in the tumor cohorts within the cerebral space is illustrated in the top right of Figure 3, revealing group-specific spatial distributions. To evaluate brain-wide structural changes, voxel-wise Euclidean distances were computed on the first PCA component, producing variability maps across groups (Figure 3, top left). The shorter-term survival group showed greater distances between pre- and post-surgery scans, suggesting more pronounced structural alterations. Moreover, this group exhibited higher spatial variability in the tumor PCA component, both before (grayscale) and after surgery (red), suggesting increased heterogeneity in tumor location and size.
@@
To quantify global structural variability, we computed voxel-wise Euclidean distances across PCA components, generating variability maps that highlight key differences between groups. This approach allowed us to capture the overall magnitude of structural differences at each voxel, revealing patterns of brain alterations associated with disease progression. To ensure a robust characterization, we evaluated both the magnitude and orientation of variations in PCA space, comparing pre- and post-surgery subgroups to assess changes relative to disease severity. These maps are displayed in Figure 3 (Global Variability Maps section) and offered a global depiction of structural variability across the brain, highlighting areas where voxel-wise differences between pre- and post-surgery scans were most pronounced in each survival group.
@@
Longer-term survivals
@@
Shorter-term survivals

### v2

@@ Abstract
Brain tumor resection is a highly complex procedure with profound implications for survival and quality of life. Predicting patient outcomes is crucial to guide clinicians in balancing oncological control with preservation of neurological function. However, building reliable prediction models is severely limited by the rarity of curated datasets that include both pre- and post-surgery imaging, given the clinical, logistical and ethical challenges of collecting such data. In this study, we develop a novel framework that integrates explainable artificial intelligence (XAI) with neuroimaging-based feature engineering for survival assessment in brain tumor patients. We curated structural MRI data from 49 patients scanned pre- and post-surgery, providing a rare resource for identifying survival-related biomarkers. A key methodological contribution is the development of a global explanation optimizer, which refines survival-related feature attribution in deep learning models, thereby improving both the interpretability and reliability of predictions. From a clinical perspective, our findings provide important evidence that survival after oncological surgery is influenced by alterations in regions related to cognitive and sensory functions. These results highlight the importance of preserving areas involved in decision-making and emotional regulation to improve long-term outcomes. From a technical perspective, the proposed optimizer advances beyond state-of-the-art XAI methods by enhancing both the fidelity and comprehensibility of model explanations, thus reinforcing trust in the recognition patterns driving survival prediction. This work demonstrates the utility of XAI-driven neuroimaging analysis in identifying survival-related variability and underscores its potential to inform precision medicine strategies in brain tumor treatment.
@@ 1 Introduction
Gliomas, the most frequent primary brain tumors, vary in aggressiveness, prognosis, and histopathology. Their treatment often involves surgical resection, followed by radiotherapy and chemotherapy. The extent of resection significantly affects survival, with surgery needing to balance tumor removal and brain function preservation [1], a principle often referred to as onco-functional balance. Beyond the immediate surgical outcome, post-operative brain reorganisation plays a central role in functional recovery. However, the mechanisms underlying these structural and functional adaptations remain insufficiently understood [2]. A more accurate characterization of these processes is essential for guiding clinical decisions, improving rehabilitation, and ultimately enhancing patient survival and quality of life.
Structural Magnetic Resonance Imaging (sMRI) provides high-resolution insights into the effects of tumor resection on brain structure, but its high dimensionality and complexity pose major analytical challenges. Machine learning (ML) techniques, particularly dimensionality reduction methods such as Principal Component Analysis (PCA) [3] or Uniform Manifold Approximation and Projection (UMAP) [4], allow for the extraction of low-dimensional representations that capture meaningful structural variations. These representations facilitate the identification of hidden patterns that may be otherwise invisible in conventional analyses. At the same time, eXplainable Artificial Intelligence (XAI) frameworks, such as feature attribution methods and model interpretability frameworks [5, 6], are increasingly recognised as essential to translate ML findings into clinically interpretable biomarkers, enabling trust and adoption in medical practice.
Despite these advances, most existing studies in brain tumour research have focused on pre-operative imaging, diagnosis, or histological classification. Much less attention has been given to post-surgical structural changes and their relationship with survival, in part due to the scarcity of longitudinal datasets covering both pre- and post-operative stages. The dataset collected and used in this study provides a rare opportunity to directly investigate these dynamics, offering insights into how surgery reshapes brain structure and how such changes relate to long-term outcomes.
In this work, we introduce a novel computational framework that combines neuroimaging-based feature engineering with a global explanation optimizer to investigate structural brain reorganization in glioma patients. The framework is designed to identify survival-related biomarkers while enhancing the stability, fidelity, and clarity of model explanations, thereby minimizing inter-method variability. Utilizing a uniquely curated dataset of pre- and post-surgery sMRI scans, we further examine how surgery-affected brain regions influence survival outcomes. Our ultimate goal is to provide clinically actionable insights that can guide surgical decision-making, refine risk stratification, and support personalized rehabilitation strategies.
The key contributions of this study are:
  • A global explanation optimizer that strengthens the reliability, fidelity, and clarity of survival-related neuroimaging biomarkers.
  • An integrated framework that combines latent-space feature engineering with XAI to provide interpretable assessments of post-surgical brain reorganization
  • Clinical insights into survival and recovery, delivering actionable guidance to neurosurgeons for optimizing surgical strategies, minimizing complications, and tailoring patient-specific post-operative care.
To the best of our knowledge, this is the first study to systematically integrate latent-space analysis of sMRI with an XAI optimization framework in the context of brain tumor surgery. By addressing both methodological and clinical challenges, our work goes beyond diagnosis to model how surgery impacts brain structure and survival. This positions our approach at the intersection of methodological innovation and clinical applicability, with direct implications for improving both acute surgical outcomes and long-term patient management.
2 Related work
Machine learning techniques have shown promising results in brain tumor analysis and outcome prediction for neurosurgical patients. Compared to conventional statistical methods, ML algorithms have demonstrated superior performance in predicting postoperative complications [7, 8] and inpatient length of stay [9]. Beyond clinical outcomes, ML-based approaches have also been widely applied to neuroimaging tasks such as brain tumor segmentation and classification [10, 11], often achieving state-of-the-art accuracy. More recent studies have started to explore pre- and post-operative MRI data, for instance to predict recovery trajectories or assess surgical effects on brain anatomy [12, 13, 14]. Similarly, ML models have been employed to predict long-term neurosurgical outcomes, including survival, recurrence, and symptom progression [15, 16]. These works highlight the increasing role of ML in neurosurgery, though the majority remain focused on diagnostic, segmentation, or histopathological classification tasks, rather than on structural reorganization after surgery.
Latent space representations, such as those derived from PCA or other manifold learning methods, have been shown to capture complex patterns in neuroimaging data that are not easily observable in raw high-dimensional spaces. They have been successfully applied in tasks ranging from correlation representation learning in multi-modal MRI segmentation [17, 18] to dimensionality reduction for group-level analyses. Such methods enable interpretable visualization and clustering of subtle neuroanatomical variations. However, to the best of our knowledge, no previous studies have leveraged latent spaces to systematically investigate longitudinal structural changes in brain tumors before and after surgery, nor their relationship with survival. This gap motivates the present study.
Alongside dimensionality reduction, the development of XAI methods has been pivotal in translating ML findings into clinically actionable knowledge [19, 20, 6, 21]. Local XAI methods provide interpretations of individual model predictions, whereas global methods offer cohort-level insights into the model’s overall decision-making process, thereby enhancing our understanding of its behavior across populations. In neuroimaging, XAI has been used to highlight relevant regions or modalities associated with tumour detection, disease progression, and prognosis, thereby increasing transparency in clinical AI. Nonetheless, a persistent challenge is the variability of explanations across methods: different local or global XAI techniques often yield inconsistent attributions, which may reduce trust in model-derived biomarkers and hinder clinical translation [6]. To the best of the authors’ knowledge, no global optimal solution exists to address inter-method variability in explanations. Bridging this gap is a key objective of our proposed work.
In summary, while ML has been widely applied to neurosurgery and brain tumor analysis, latent space methods remain underexplored for modelling structural reorganization, and current XAI techniques lack robustness when applied to survival-related neuroimaging biomarkers. Our work bridges these gaps by (i) introducing a latent-PCA framework to capture post-surgical structural changes in glioma patients, and (ii) proposing a global explanation optimizer to mitigate inter-method variability in XAI, thereby offering more stable and clinically interpretable biomarkers of survival.
3 Methods
@@
This work combines latent space feature engineering and XAI methods to identify biomarkers related to surgical outcomes. A summary of the implemented framework is presented in Fig. 1. The main part consists of two phases: feature engineering through dimensionality reduction and a global explanation optimizer integrated with DL networks and XAI methods.
@@ 3.1 Phase I: Feature engineering based on dimensionality reduction
We utilized PCA to extract the most relevant patterns of variation across the brain and tumor cohorts, considering four groups defined by time (pre- vs. post-surgery) and survival (longer-term vs. shorter-term). Two main approaches were used: first, analyzing PCA component variability across groups, and second, quantifying variability across PCA components (see Fig. 1):
@@ 3.1.2 Variability quantification across PCA Components
3.2 Phase II: Feature identification based on cohort-level explanations, integrated with DL networks and tailored to survival classification
Fig. 1 illustrates the explainable AI framework developed to identify global (cohort-level) patterns associated with survival outcomes following brain tumor surgery. Given the limited size of our clinical dataset and the need to avoid overfitting, we first trained a generalized unsupervised model on a large, heterogeneous dataset of structural MRI brain tumor scans. This encoder–decoder architecture learned the distribution of sMRI data, capturing variability and heterogeneity across patients to reduce bias in downstream analyses. Building on this foundation, a binary classification model was trained and validated to distinguish patients with shorter versus longer survival. To optimize performance, we systematically evaluated different strategies, including freezing versus fine-tuning encoder layers and conducting an ablation study of alternative network architectures derived from the unsupervised stage. Cohort-level explanations were then integrated into the survival classification task using our proposed global explanation optimizer, which enhanced both the clarity and the consistency of global survival-related patterns.
@@ 3.2.1 Unsupervised learning of structural MRI
Two deep learning architectures were employed in the unsupervised learning stage: a convolutional autoencoder (AE) with three encoder and three decoder blocks, and the Swin-Unet [22]. Both models were trained to reconstruct full pre- and post-operative 3D structural MRI scans in an unsupervised setting. To assess reconstruction performance and generalization capacity, we performed an ablation study comparing two different cohort training strategies. Further implementation details are provided in Section 4.2.
@@ 3.2.4 The proposed global explanation optimizer of structural MRI
We extracted the first three principal components via PCA from the total cohort saliency maps, generated using three of the six widely used attribution methods employed in this study: Guided Backpropagation, Guided Grad-CAM, and Gradient SHAP. These components, along with their weighted average, calculated according to the procedure described in [6], were used as four inputs to the proposed global explanation optimizer.
@@
3.3 Summary of the proposed framework
The proposed framework provides a unified pipeline to analyze post-surgical brain structural changes and identify survival-related biomarkers from MRI data. As shown in Fig. 1, the workflow has a clear flow from raw imaging data to clinically interpretable outcomes. Before the main analysis, all sMRI images are preprocessed to ensure consistency across subjects. This includes spatial alignment, skull stripping, and masking of tumor regions. These steps harmonize the images and reduce variability unrelated to brain structure. The framework then proceeds in two complementary phases, combining latent-space feature engineering and explainable AI methods to extract meaningful patterns and biomarkers:
  1. 1.
    Phase I – Latent-Space Feature Engineering: Pre- and post-surgery MRI scans are transformed into low-dimensional latent spaces using PCA. This step captures the most relevant patterns of variability across the brain and tumor cohorts, grouped by time (pre- vs. post-surgery) and survival (longer-term vs. shorter-term). By quantifying both local and global variability across PCA components and groups, this phase provides a comprehensive view of structural changes induced by surgery.
  2. 2.
    Phase II – Cohort-Level Feature Identification via XAI: Latent representations are then used to train survival classifiers based on DL encoders. Multiple local XAI techniques are applied to the trained models to produce individual-level explanations. These explanations are combined and optimized using our global explanation optimizer, yielding stable and interpretable cohort-level maps of brain regions associated with survival outcomes.
The framework therefore bridges unsupervised feature extraction and explainable deep learning, enabling the identification of meaningful structural patterns while ensuring robustness and interpretability. The final outputs are global explanation maps, highlighting key brain regions and tumor areas linked to survival.
4 Experimental settings
@@ 4.1 Dataset
[…]
In Phase II, an additional dataset from the 2025 Brain Tumor Segmentation (BraTS) Glioma Challenge [32] was employed. Hereafter, we refer to this dataset as BraTS2025. This dataset comprises pre- and post-treatment T2-weighted MRI scans. We used a total of 1453 images (1251 pre-treatment and 202 post-treatment). These scans were used to train the unsupervised learning models (see 3.2.1). Demographic information was not provided for this dataset.
@@ 4.2 Implementation details
For survival classification, sparse categorical cross-entropy was used as the loss function, optimized with Adam. The learning rate was constant for the first 100 epochs and then reduced by a factor of 0.1 every 100 epochs. Early stopping was applied after 100 epochs of no improvement (max 400 epochs). A 5-fold CV was used. Both tasks employed data augmentation, including rotation ($15^{\circ},15^{\circ}]$), width/height shift (up to 20 pixels), and intensity shift (up to 20%). Hyperparameter tuning tested learning rates: $5\times 10^{-2}$, $5\times 10^{-3}$, $5\times 10^{-4}$, and $5\times 10^{-5}$ (see Fig. 2a.). The XAI task used the Adam optimizer, but no data augmentation. The cost function was (4). For 3D tasks, training lasted up to 100 epochs, with early stopping after 10 epochs of no improvement beyond the first 50. Hyperparameter tuning tested the same learning rates as previously and various combinations of the $l_{1}$, $l_{2}$, and $l_{3}$ parameters in (4) with the best combination of parameters determined as $l_{1}=0.4$, $l_{2}=0.3$, $l_{3}=0.3$ and a learning rate $5\times 10^{-5}$ (see Fig. 2b.). Codes were implemented in Python using PyTorch and trained on one A100 GPU with 64 GB RAM. It will be publicly available on GitHub.
@@ 4.3.2 Sparseness Metric
[…]
@@ 5.1 Structural patterns identified using feature engineering based on PCA
Once the PCA components (brain: 8, tumor: 12) were computed across groups, structural variability was quantified to explore spatial differences in tumor and brain patterns. The localization of the first PCA component in the tumor cohorts within the cerebral space is illustrated in the top right of Fig.3, revealing group-specific spatial distributions. To evaluate brain-wide structural changes, voxel-wise Euclidean distances were computed on the first PCA component, producing variability maps across groups (Fig.3, top left). The shorter-term survival group showed greater distances between pre- and post-surgery scans, suggesting more pronounced structural alterations. Moreover, this group exhibited higher spatial variability in the tumor PCA component, both before (grayscale) and after surgery (red), suggesting increased heterogeneity in tumor location and size.
@@
To quantify global structural variability, we computed voxel-wise Euclidean distances across PCA components, generating variability maps that highlight key differences between groups. This approach allowed us to capture the overall magnitude of structural differences at each voxel, revealing patterns of brain alterations associated with disease progression. To ensure a robust characterization, we evaluated both the magnitude and orientation of variations in PCA space, comparing pre- and post-surgery subgroups to assess changes relative to disease severity. These maps are displayed in Fig. 3 (Global Variability Maps section) and offered a global depiction of structural variability across the brain, highlighting areas where voxel-wise differences between pre- and post-surgery scans were most pronounced in each survival group.
@@
Longer-term
survivals
@@
Shorter-term
survivals